A,B and C can do a piece of work together in 3 days. A can do it alone in 10 days and B can do it alone in 8 days. How long will it take C to do the job alone?: A,B and C can do a piece of work together in 3 days. A can do it alone in 10 days and B can do it alone in 8 days. How long will it take C to do the job alone?
Answer by edjones(2401)

A can do 1/10 of the job in one day.
B can do 1/8 of the job in one day.
.
1/10 + 1/8 + c = 1/3
12/120 + 15/120 + c = 40/120
c=40/12 - 27/120
=13/120 of the job in one day
So c can do the job in 120/13 days or 9 3/13 days.
.
Ed